OpenStudy (anonymous):

Hans pays $205 in advance on his account at the athletic club. Each time he uses the club, $10 is deducted from the account. Write an equation that represents the value remaining in his account after x visits to the club. Find the value remaining in the account after 11 visits.

OpenStudy (anonymous):

y = 205 - 10(11) = 95 im guessing

OpenStudy (anonymous):

\[205 - 10x \] \[205 - 10 * 11 \] \[205 - 110 = 95\]

OpenStudy (anonymous):

indeed, thank you

OpenStudy (anonymous):

no problem :)